25angianok:

A table contains the ordered pairs (3, 42.25) and (5, 50.75). If the relationship in the table is linear, explain how to find the initial value.

mxddi3:

sorry no one got to you sooner :/ here's something to hopefully help you: so we have (3, 42.25) and (5, 50.75) The linear relationship ( the linear equation ): y = m x + b ( the slope-intercept form ) if m is the slope, we need to do \[\frac{ y2-y1 }{ x2-x1 }\] m = ( 50.75 - 42.25 ) / ( 5 - 3 ) = 8.50 / 2 = 4.25 the slope is 4.25. now, using the first point fill in your y,m,and x: 42.25 = 4.25 * 3 + b 42.25=12.75+b then, : b = 42.25 - 12.75 b = 29.50 y = 4.25 * x + 29.50 Answer: The initial value is : ( 0, 29.50 ) or 29.50.